(n to the 10th power) (n negative to the 10th power) How to I foil this?
Answer by stanbon(75887)

(n to the 10th power) (n negative to the 10th power) How to I foil this?
-------------
FOIL does not apply to your problem because the factors are not binomials.
----------------
(n^10)(n^-10) = n^0 = 1 as long as n does not equal zero.
----------------
Cheers,
Stan H.